If you want a poll to have a margin of error of 2.43%, how large will your sample have to be? Round your answer to the nearest whole number.
Answer #1

The sample size formula is given as below:

n = p*q*(Z/E)^2

We are not given the estimate for proportion, so we take

p = 0.5

q = 1 – p = 0.5

WE assume Confidence level = 95%

Critical Z value = 1.96

(by using z-table)

Margin of error = E = 0.0243

The sample size is given as below:

n = p*q*(Z/E)^2

n = 0.5*0.5*(1.96/0.0243)^2

n = 1626.446

Required sample size = 1627